### v1

# Equality in Liakopoulos’s generalized dual Loomis–Whitney inequality via Barthe’s Reverse Brascamp–Lieb inequality

Abstract.
We use the characterization of the case of equality in Barthe’s Geometric Reverse Brascamp–Lieb inequality in order to characterize equality in Liakopoulos’s volume estimate in terms of sections by certain lower-dimensional linear subspaces.
1. Introduction and results
Geometric inequalities, where the volume of a shape is given and we search the extremal value of a quantity, have been considered since the time of ancient Greeks and have been at the center of attention in Geometry and Analysis in the last two centuries. The most classical one is probably the Isoperimetric inequality. Let $B^{n}\subset\mathbb{R}^{n}$ be the origin-centered Euclidean unit ball, and let $X\subset\mathbb{R}^{n}$ be a compact set such that ${\rm int}\,X\neq\emptyset$ and its boundary $\partial X$ is an $(n-1)$-dimensional Lipschitz manifold. Let $|X|$ denote the Lebesgue measure of $X$. We define the surface area $S(X)$ of $X$ as $S(X)=\lim_{r\to 0^{+}}\frac{|X+rB^{n}|-|X|}{r}$. The Isoperimetric inequality states that
$S(X)\geq n|B^{n}|^{\frac{1}{n}}\cdot|X|^{\frac{n-1}{n}},$
with equality if and only if $X$ is a Euclidean ball. One of the major results of Bent Fuglede was that he proved the optimal stability version of the Isoperimetric inequality for star-shaped $X\subset\mathbb{R}^{n}$ with Lipschitz boundary in terms of the Hausdorff distance in a series of papers [34, 35, 36] in 1986–1991 (see Figalli, Maggi, Pratelli [33] and Fusco, Maggi, Pratelli [37] for optimal stability results in terms of volume difference). Here the Isoperimetric inequality is a consequence of the Brunn–Minkowski inequality
$|\alpha X+\beta Y|^{\frac{1}{n}}\geq\alpha|X|^{\frac{1}{n}}+|Y|^{\frac{1}{n}}$
for compact sets $X,Y\subset\mathbb{R}^{n}$ with $|X|,|Y|>0$ and $\alpha,\beta>0$, where equality holds if and only if $X$ and $Y$ are convex and homothetic; namely, $Y=\lambda X+z$ for some $\lambda>0$ and $z\in\mathbb{R}^{n}$. See Figalli, Maggi, Pratelli [33] for an elegant proof of the Brunn–Minkowski inequality including equality conditions and even a stability version based on optimal transport. In turn, the Brunn–Minkowski inequality has a functional form that is called Prékopa–Leindler inequality, which we recall in a form most suitable for use in this paper: If $\sum_{i=1}^{k}c_{i}=1$ holds for $c_{1},\ldots,c_{k}>0$ with $k\geq 2$, then for any non-negative $f_{i}\in L_{1}(\mathbb{R}^{n})$, $i=1,\ldots,k$, we have
$\int_{\mathbb{R}^{n}}^{*}\sup_{x=\sum_{i=1}^{k}c_{i}x_{i}}\;\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,dx\geq\prod_{i=1}^{k}\left(\int_{\mathbb{R}^{n}}f_{i}\right)^{c_{i}}$ (1)
where $\int_{\mathbb{R}^{n}}^{*}$ stands for outer integral as the integrand on the left hand side may not be measurable (see Gardner [38] for a survey on various aspects of the Brunn–Minkowski inequality and the Prékopa–Leindler inequality). If $\int_{\mathbb{R}^{n}}f_{i}>0$ for all $i=1,,\ldots,k$, then equality in the Prékopa–Leindler inequality (1) implies that there exist $a_{1},\ldots,a_{k}>0$, $w_{1},\ldots,w_{k}\in\mathbb{R}^{n}$ and a log-concave function $\varphi$ such that $\sum_{i=1}^{k}c_{i}w_{i}=o$ and $f_{i}(x)=a_{i}\varphi(x+w_{i})$ for a.e. $x\in\mathbb{R}^{n}$ (see Dubuc [30] and Balogh, Kristály [6] for the equality case, and Figalli, van Hintum, Tiba [32] for a stability version). Here a function $f:\,\mathbb{R}^{n}\to[0,\infty)$ is log-concave if $f((1-\lambda)x+\lambda y)\geq f(x)^{1-\lambda}f(y)^{\lambda}$ holds for $x,y\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$; or equivalently, $f=e^{-\varphi}$ for a convex function $\varphi:\,\mathbb{R}^{n}\to(-\infty,\infty]$.
In this note, we characterize equality in Liakopoulos’s inequality (Theorem 5), as a first step towards a possible stability version. However, the story of this paper originates from the classical Loomis–Whitney inequality (2) proved in [46].
Let $e_{1},\ldots,e_{n}$ denote an orthonormal basis of $\mathbb{R}^{n}$. For a compact set $K\subset\mathbb{R}^{n}$ with ${\rm dim}\,{\rm aff}\,K=m$, we write $|K|$ for the $m$-dimensional Lebesgue measure of $K$. For a proper linear subspace $E$ of $\mathbb{R}^{n}$ ($E\neq\mathbb{R}^{n}$ and $E\neq\{0\}$), let $P_{E}$ denote the orthogonal projection into $E$.
Theorem 1 (Loomis, Whitney).
If $K\subset\mathbb{R}^{n}$ is compact and $|K|>0$, then
$|K|^{n-1}\leq\prod_{i=1}^{n}|P_{e_{i}^{\bot}}K|,$ (2)
with equality if and only if $K=\oplus_{i=1}^{n}K_{i}$, where $\mathrm{aff}\,K_{i}$ is a line parallel to $e_{i}$.
In order to consider a generalization of the Loomis–Whitney inequality, we set $[n]:=\{1,\ldots,n\}$, and for a non-empty proper subset $\sigma\subset[n]$, we define $E_{\sigma}=\operatorname{lin}\{e_{i}\}_{i\in\sigma}$. For $s\geq 1$, we say that the not necessarily distinct proper non-empty subsets $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ if each $j\in[n]$ is contained in exactly $s$ of $\sigma_{1},\ldots,\sigma_{k}$. The Bollobás–Thomason inequality [18] is as follows.
Theorem 2 (Bollobás, Thomason).
If $K\subset\mathbb{R}^{n}$ is compact, and $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ for an integer $s\geq 1$, then
$|K|^{s}\leq\prod_{i=1}^{k}|P_{E_{\sigma_{i}}}K|.$ (3)
Remarks.
  • (i)
    The case when $k=n$ and $s=n-1$, and so when we may assume that $\sigma_{i}=[n]\backslash\{i\}$, is the Loomis–Whitney inequality Theorem 1.
  • (ii)
    According to Bennett, Carbery, Christ, Tao [17], a collection of linear subspaces $E_{1},\ldots,E_{k}\subset\mathbb{R}^{n}$ , each of dimension at least one, together with positive constants $c_{1},\ldots,c_{k}>0$ forms a Geometric Brascamp–Lieb datum if
    $\sum_{i=1}^{k}c_{i}P_{E_{i}}=I_{n},$ (4)
    where $I_{n}$ is the identity map on $\mathbb{R}^{n}$. The term ”geometric” refers to the fact that the surjective linear maps $P_{E_{i}}\colon\mathbb{R}^{n}\to E_{i}$, $i=1,\ldots,k$ are projections.
    Now the the subspaces $E_{\sigma_{i}}:={\rm lin}\{e_{j}:j\in\sigma_{i}\}$ in Theorem 2 satisfy
    $\sum_{i=1}^{k}\frac{1}{s}\,P_{E_{\sigma_{i}}}=I_{n};$ (5)
    that is, they form a Geometric Brascamp–Lieb datum, where each coefficients is $\frac{1}{s}$.
The equality case of the Bollobás–Thomason inequality Theorem 2, based on Valdimarsson [57], has been known to the experts. Let $s\geq 1$, and let $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ be an $s$-uniform cover of $[n]$. We say that the disjoint union $[n]=\tilde{\sigma}_{1}\sqcup\ldots\sqcup\tilde{\sigma}_{\ell}$ forms the partition of $[n]$ induced by the $s$-uniform cover $\sigma_{1},\ldots,\sigma_{k}$ if $\{\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{\ell}\}$ consists of all non-empty distinct subsets of $[n]$ of the form $\cap_{i=1}^{k}\sigma_{i,0}$, where for each $i$, either $\sigma_{i,0}=\sigma_{i}$ or $\sigma_{i,0}=[n]\setminus\sigma_{i}$.
Theorem 3 (Folklore).
Let $K\subset\mathbb{R}^{n}$ be compact with $|K|>0$, and let $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ for an integer $s\in\{1,\ldots,n-1\}$. Then equality holds in (3) if and only if $K=\oplus_{i=1}^{\ell}P_{E_{\tilde{\sigma}_{i}}}K$, where $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{\ell}$ form the partition of $[n]$ induced by $\sigma_{1},\ldots,\sigma_{k}$.
We note that Ellis, Friedgut, Kindler, Yehudayoff [29] even proved a stability version of the Bollobás–Thomason inequality Theorem 2 in the case relevant to information theory, and hence that of the Loomis–Whitney inequality, as well.
The real starting point of our note is Meyer’s dual Loomis–Whitney inequality in [53], where equality holds for affine cross-polytopes.
Theorem 4 (Meyer).
If $K\subset\mathbb{R}^{n}$ is a compact convex set with $o\in\mathrm{int}\,K$, then
$|K|^{n-1}\geq\frac{n!}{n^{n}}\prod_{i=1}^{n}\left|K\cap e_{i}^{\bot}\right|,$ (6)
with equality if and only if $K=\operatorname{conv}\{\pm\lambda_{i}e_{i}\}_{i=1}^{n}$ for $\lambda_{i}>0$, $i=1,\ldots,n$.
Let us observe that unlike in the case of the Loomis-Whitney inequality, the assumption that $K$ is convex is necessary in Meyer’s inequality (6). Without convexity, $|K|$ can be small even if all sections $K\cap e_{i}^{\perp}$ have large volume.
We note that various reverse and dual Loomis–Whitney type inequalities were proved by Campi, Gardner, Gronchi [48], Brazitikos et al. [22, 23], Alonso-Gutiérrez et al. [1, 2]. In particular, according to (5), the following theorem, due to Liakopoulos [43], is a generalization of Meyer’s inequality on the one hand, and is the generalized dual form of the Bollobás–Thomason inequality Theorem 2 on the other hand.
Theorem 5 (Liakopoulos).
If $K\subset\mathbb{R}^{n}$ is a compact convex set with $o\in\operatorname{int}K$, and the linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$, with ${\rm dim}\,E_{i}=d_{i}\in\{1,\ldots,n-1\}$, form a Geometric Brascamp–Lieb datum; that is,
$\sum_{i=1}^{k}c_{i}P_{E_{i}}=I_{n},$ (7)
then
$|K|\geq\frac{\prod_{i=1}^{k}d_{i}!}{n!}\cdot\prod_{i=1}^{k}\left|K\cap E_{i}\right|^{c_{i}}.$ (8)
Like in Meyer’s inequality, the convexity of $K$ is a necessary condition in Liakopoulos’s inequality.
Given the linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ as part of a Geometric Brascamp–Lieb datum in Theorem 5, we say that $F=\cap_{i=1}^{k}E_{i,0}$ is an independent subspace if either $E_{i,0}=E_{i}$ or $E_{i,0}=E_{i}^{\bot}$ for any $i=1,\ldots,k$, and ${\rm dim}\,F\geq 1$. In general, there may not exist any independent subspace associated to $E_{1},\ldots,E_{k}$; however, if $F\neq F^{\prime}$ are independent subspaces, then $F^{\prime}\subset F^{\bot}$. Our main result characterizes equality in Liakopoulos’s Theorem 5.
Theorem 6.
We have equality in (8) in Theorem 5 if and only if
  • (i)
    $\oplus_{j=1}^{\ell}F_{j}=\mathbb{R}^{n}$ holds for the independent subspaces $F_{1},\ldots,F_{\ell}\subset\mathbb{R}^{n}$ associated to $E_{1},\ldots,E_{k}$, and
  • (ii)
    $K=\operatorname{conv}\{K\cap F_{j}\}_{j=1}^{\ell}$.
Concerning the structure of the paper: we introduce the Brascamp–Lieb inequality and Barthe’s reverse form, which has a crucial role in the proof of Theorem 6, in Section 2. Finally, Theorem 6 is verified in Section 3.
2. Equality in Barthe’s Geometric Reverse Brascamp–Lieb ineuqlaity and some related results
For a non-trivial linear subspace $E\subset\mathbb{R}^{n}$ ($E\neq\{o\}$), we write $P_{E}$ to denote the orthogonal projection into $E$ where $P_{E}=I_{n}$ is the identity if $E=\mathbb{R}^{n}$. In addition, the linear subspaces $E_{1},\ldots,E_{k}\subset\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$ form a Geometric Brascamp–Lieb datum if they satisfy
$\sum_{i=1}^{k}c_{i}P_{E_{i}}=I_{n}.$ (9)
Comparing traces in (9) implies that if $n_{i}=\dim E_{i}$, then
$\sum_{i=1}^{k}c_{i}n_{i}=n.$ (10)
The term “Geometric Brascamp–Lieb datum”, introduced by Bennett, Carbery, Christ, Tao in [17], was motivated by Theorem 7. Theorem 7 originates in the works of Brascamp, Lieb [21] and Ball [3, 4] in the rank one case, ($\dim E_{i}=1$ for $i=1,\ldots,k$), and Lieb [47] and Barthe [8] in the general case. In the rank one case, the Geometric Brascamp–Lieb datum is known as a Parseval frame in coding theory and computer science (see, for example, Casazza, Tran, Tremain [26]).
Theorem 7 (Brascamp, Lieb, Ball, Barthe).
For the non-trivial linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and constants $c_{1},\ldots,c_{k}>0$ satisfying (9), and for non-negative $f_{i}\in L_{1}(E_{i})$, we have
$\int_{\mathbb{R}^{n}}\prod_{i=1}^{k}f_{i}(P_{E_{i}}x)^{c_{i}}\,dx\leq\prod_{i=1}^{k}\left(\int_{E_{i}}f_{i}\right)^{c_{i}}$ (11)
Remark 8.
This is Hölder’s inequality if $E_{1}=\ldots=E_{k}=\mathbb{R}^{n}$, and hence $\sum_{i=1}^{k}c_{i}=1$ (cf. (10)).
According to (5), the Geometric Brascamp–Lieb inequality (Theorem 7) directly yields the Bollobás–Thomason inequality (Theorem 5) by choosing $E_{i}=E_{\sigma_{i}}$, $c_{i}=\frac{1}{s}$ and $f_{i}=\mathbf{1}_{P_{E_{\sigma_{i}}}}$ for $i=1,\ldots,k$.
We note that equality holds in Theorem 7 if $f_{i}(x)=e^{-\pi\|x\|^{2}}$ for $i=1,\ldots,k$, that is, when each $f_{i}$ is a Gaussian density function. Theorem 7 is an important special case of the general Brascamp–Lieb inequality Theorem 14, discovered by Ball [4, 5] in the rank one case and by Barthe [8] in the general case.
After partial results by Barthe [8], Carlen, Lieb, Loss [25] and Bennett, Carbery, Christ, Tao [17], it was Valdimarsson [57] who characterized equality in the Geometric Brascamp–Lieb inequality. In order to state his result, we need some notation. Let the proper linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and constants $c_{1},\ldots,c_{k}>0$ satisfy (9). In order to understand extremizers in (14), following Carlen, Lieb, Loss [25] and Bennett, Carbery, Christ, Tao [17], we say that a non-zero linear subspace $V$ is a critical subspace if
$\sum_{i=1}^{k}c_{i}\dim(E_{i}\cap V)=\dim V,$
which is, in turn, equivalent to
$E_{i}=(E_{i}\cap V)+(E_{i}\cap V^{\bot})$ for $i=1,\ldots,k$
according to [17].
Valdimarsson [57] introduced the so-called independent subspaces and the dependent space. We write $J$ to denote the set of $2^{k}$ functions $\{1,\ldots,k\}\to\{0,1\}$. If $\varepsilon\in J$, then let $F_{(\varepsilon)}=\cap_{i=1}^{k}E_{i}^{(\varepsilon(i))}$, where $E_{i}^{(0)}=E_{i}$ and $E_{i}^{(1)}=E_{i}^{\bot}$ for $i=1,\ldots,k$. Let $J_{0}$ denote the subset of $J$ such that $\dim F_{(\varepsilon)}\geq 1$ for all $\varepsilon\in J_{0}$. Such an $F_{(\varepsilon)}$ is called independent following Valdimarsson [57]. Readily, $F_{(\varepsilon)}$ and $F_{(\tilde{\varepsilon})}$ are orthogonal if $\varepsilon\neq\tilde{\varepsilon}$ for $\varepsilon,\tilde{\varepsilon}\in J_{0}$. In addition, we write $F_{\mathrm{dep}}$ to denote the orthogonal complement of $\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}$ in $\mathbb{R}^{n}$. Thus, $\mathbb{R}^{n}$ decomposes into a direct sum of pairwise orthogonal linear subspaces in the form
$\mathbb{R}^{n}=\left(\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}\right)\oplus F_{\mathrm{dep}}.$ (12)
In (12), it is possible that $J_{0}=\emptyset$, and so $\mathbb{R}^{n}=F_{\mathrm{dep}}$, or $F_{\mathrm{dep}}=\{0\}$, and hence $\mathbb{R}^{n}=\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}$ in that case.
For a non-zero linear subspace $L\subset\mathbb{R}^{n}$, we say that a linear transformation $A:\,L\to L$ is positive definite if $\langle Ax,y\rangle=\langle x,Ay\rangle$ and $\langle x,Ax\rangle>0$ for any $x,y\in L\backslash\{0\}$.
Theorem 9 (Valdimarsson).
For the proper linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$ satisfying (9), let us assume that equality holds in the Brascamp–Lieb inequality (11) for non-negative $f_{i}\in L_{1}(E_{i})$, $i=1,\ldots,k$. If $F_{\mathrm{dep}}\neq\mathbb{R}^{n}$, then let $F_{1},\ldots,F_{\ell}$ be the independent subspaces, and if $F_{\mathrm{dep}}=\mathbb{R}^{n}$, then let $\ell=1$ and $F_{1}=\{0\}$. Then there exist $b\in F_{\mathrm{dep}}$ and $\theta_{i}>0$ for $i=1,\ldots,k$, integrable non-negative $h_{j}:\,F_{j}\to[0,\infty)$ for $j=1,\ldots,\ell$, and a positive definite matrix $A:F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ such that the eigenspaces of $A$ are critical subspaces and
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x-b\rangle}\prod_{F_{j}\subset E_{i}}h_{j}(P_{F_{j}}(x))\mbox{ \ \ \ for Lebesgue a.e. $x\in E_{i}$}.$ (13)
On the other hand, if for any $i=1,\ldots,k$, $f_{i}$ is of the form as in (13), then equality holds in (11) for $f_{1},\ldots,f_{k}$.
Theorem 9 explains the term ”independent subspaces” because the functions $h_{j}$ on $F_{j}$ are chosen freely and independently from each other.
A reverse form of the Geometric Brascamp–Lieb inequality was proved by Barthe [8]. We write $\int^{*}_{\mathbb{R}^{n}}\varphi$ to denote the outer integral for a possibly non-integrable function $\varphi:\,\mathbb{R}^{n}\to[0,\infty)$; namely, the infimum (actually, minimum) of $\int_{\mathbb{R}^{n}}\psi$ where $\psi\geq\varphi$ is Lebesgue measurable.
Theorem 10 (Barthe).
For the non-trivial linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and $c_{1},\ldots,c_{k}>0$ satisfying (9), and for non-negative $f_{i}\in L_{1}(E_{i})$, we have
$\int_{\mathbb{R}^{n}}^{*}\sup_{x=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\;\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,dx\geq\prod_{i=1}^{k}\left(\int_{E_{i}}f_{i}\right)^{c_{i}}.$ (14)
Remark 11.
This is the Prékopa–Leindler inequality (1) if $E_{1}=\ldots=E_{k}=\mathbb{R}^{n}$ then $P_{E_{i}}=I_{n}$, and hence $\sum_{i=1}^{k}c_{i}=1$ (cf. (10)).
We note that Barthe’s Geometric Reverse Brascamp–Lieb inequality (Theorem 10) directly yields Liakopoulos’s inequality in Theorem 5 by choosing $f_{i}=\mathbf{1}_{K\cap E_{i}}$ for $i=1,\ldots,k$.
We say that a function $h\colon\mathbb{R}^{n}\to[0,\infty)$ is log-concave if $h((1-\lambda)x+\lambda\,y)\geq h(x)^{1-\lambda}h(y)^{\lambda}$ for any $x,y\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$; or in other words, $h=e^{-W}$ for a convex function $W\colon\mathbb{R}^{n}\to(-\infty,\infty]$. It follows that $\operatorname{supp}h=\operatorname{cl}\{x\in\mathbb{R}^{n}\colon h(x)>0\}$ is convex, it has non-empty interior if $\int_{\mathbb{R}^{n}}h>0$, and $h$ is continuous on $\operatorname{int}\operatorname{supp}h$. In addition, if $0<\int_{\mathbb{R}^{n}}h<\infty$, then $h$ is bounded, and for any $t$ with $0<t<\sup h$, the open set $\{h>t\}$ is convex and bounded. Böröczky, Kalantzopoulos, Xi [20] proved the following characterization of equality in the Barthe’s inequality (14).
Theorem 12 ([20]).
For linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and $c_{1},\ldots,c_{k}>0$ satisfying (9), if $F_{\mathrm{dep}}\neq\mathbb{R}^{n}$, then let $F_{1},\ldots,F_{\ell}$ be the independent subspaces, and if $F_{\mathrm{dep}}=\mathbb{R}^{n}$, then let $\ell=1$ and $F_{1}=\{0\}$.
If equality holds in Barthe’s Geometric Reverse Brascamp–Lieb inequality (14) for non-negative $f_{i}\in L_{1}(E_{i})$ with $\int_{E_{i}}f_{i}>0$, $i=1,\ldots,k$, then
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x-b_{i}\rangle}\prod_{F_{j}\subset E_{i}}h_{j}\left(P_{F_{j}}(x-w_{i})\right)$ (15)
for Lebesgue a.e. $x\in E_{i}$, where
  • $\theta_{i}>0$, $b_{i}\in E_{i}\cap F_{\rm dep}$ and $w_{i}\in E_{i}$ for $i=1,\ldots,k$,
  • $h_{j}\in L_{1}(F_{j})$ is non-negative for $j=1,\ldots,\ell$, and, in addition, $h_{j}$ is log-concave if there exist $\alpha\neq\beta$ with $F_{j}\subset E_{\alpha}\cap E_{\beta}$,
  • $A:F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ is a positive definite matrix such that the eigenspaces of $A$ are critical subspaces.
On the other hand, if for any $i=1,\ldots,k$, $f_{i}$ is of the form as in (15) and equality holds for all $x\in E_{i}$ in (15), then equality holds in (14) for $f_{1},\ldots,f_{k}$.
We discuss some special cases of Theorem 12 that will be useful in the proof of Theorem 6. In Lemma 13 we use the notations of Theorem 12.
Lemma 13.
Let $f_{1},\ldots,f_{k}$ form a set of extremizers of Barthe’s Geometric Reverse Brascamp–Lieb inequality.
  • (i)
    If for each $\alpha=1,\ldots,k$, the remaining subspaces $\{E_{i}\}_{i\neq\alpha}$ span $\mathbb{R}^{n}$, then each $f_{i}$ is $\mathcal{H}^{n_{i}}$ a.e. equal to some log-concave function on $E_{i}$.
  • (ii)
    If each $f_{i}$ is log-concave, then one may assume that each $h_{j}$, $j=1,\ldots,\ell$, is log-concave, and hence (15) holds for all $x\in\mathrm{relint}\,\mathrm{supp}\,f_{i}$.
  • (iii)
    If each $f_{i}$ is even, then one may assume that each $b_{i}=w_{i}=o$ for $i=1,\ldots,k$, and each $h_{j}$ is even for $j=1,\ldots,\ell$ (and actually each $h_{j}$ being even log-concave can be assumed if each $f_{i}$ is even log-concave).
Proof.
For (i), the condition that for any $\alpha=1,\ldots,k$, $\{E_{i}\}_{i\neq\alpha}$ spans $\mathbb{R}^{n}$ in Theorem 12 is equivalent to saying that for any $F_{j}$, $j=1,\ldots,\ell$, there exist $\alpha\neq\beta$ with $F_{j}\subset E_{\alpha}\cap E_{\beta}$.
In the argument for (ii) and (iii), if $F_{\mathrm{dep}}\neq\{o\}$, then we write $h_{0,i}(y)=e^{-\langle Ay,y-b_{i}\rangle}$ for $i=1,\ldots,k$ and $y\in E_{i}\cap F_{\mathrm{dep}}$, and if $F_{j}\subset E_{i}$ is an independent subspace, then we write $h_{j,i}\left(y\right)=h_{j}\left(y-P_{F_{j}}w_{i}\right)$ for $y\in F_{j}$.
For (ii), first we show that even if some independent subspace $F_{m}$ is contained in a single $E_{\alpha}$, then $h_{m}$ Lebesgue a.e. equals to some log-concave function on $F_{m}$. In this case, let $z\in E_{\alpha}\cap F_{m}^{\bot}$ such that $h_{j,\alpha}(P_{F_{j}}z)>0$ whenever $F_{j}\subset E_{\alpha}$ and $j\neq m$, and for Lebesgue a.e. $x\in z+F_{m}$, we have
$f_{\alpha}(x)=\theta_{\alpha}h_{0,\alpha}\left(P_{F_{\rm dep}}z\right)h_{m,\alpha}\left(P_{F_{m}}x\right)\prod_{F_{j}\subset E_{\alpha},\&\,j\neq m}h_{j,\alpha}\left(P_{F_{j}}z\right).$
As $f_{\alpha}$ is log-concave, $h_{m,\alpha}$ agrees with a log-concave function for Lebesgue a.e. $y\in F_{m}$, and hence the same holds for $h_{\alpha}$.
Now as log-concave functions are continuous on their support, we may assume that each $h_{j}$ is actually log-concave, and hence (15) holds for all $x\in\operatorname{relint}\operatorname{supp}f_{i}$ for all $i=1,\ldots,k$.
To show (iii), one replaces each $h_{j,i}$ by the function $y\mapsto\sqrt{h_{j,i}(y)\cdot h_{j,i}(-y)}$ that is log-concave if $h_{j,i}$ is log-concave. ∎
As it is stated in (i) of Lemma 13, if for any $\alpha=1,\ldots,k$, $\{E_{i}\}_{i\neq\alpha}$ spans $\mathbb{R}^{n}$ in Theorem 12, then the $h_{j}$ in the extremizers of the Geometric Barthe’s inequality are log-concave. The explanation for this phenomenon is the following. If $F_{j}$ is an independent subspace with $j\in\{1,\ldots,\ell\}$, then $\sum_{E_{i}\supset F_{j}}c_{i}=1$. If $f_{1},\ldots,f_{k}$ are of the form (15), then equality in Barthe’s inequality (14) yields
[…]
Therefore, if there exist $\alpha\neq\beta$ with $F_{j}\subset E_{\alpha}\cap E_{\beta}$, then the equality conditions in the Prékopa-Leindler inequality imply that each $h_{j}$ is log-concave. On the other hand, if there exists $\alpha\in\{1,\ldots,k\}$ such that $F_{j}\subset E_{\beta}^{\bot}$ for $\beta\neq\alpha$, then we do not have any condition on $h_{j}$, and $c_{\alpha}=1$.
For completeness, let us state and discuss the general Brascamp–Lieb inequality and its reverse form due to Barthe. The following theorem was proved by Brascamp, Lieb [21] in the rank one case, and Lieb [47] in the general case.
We define the Brascamp–Lieb datum in the general case. Let $H_{i}$ be $n_{i}$-dimensional Euclidean spaces, with $n_{i}\geq 1$, for $i=1,\ldots,k$. Let $B_{i}:\mathbb{R}^{n}\to H_{i}$ be surjective linear maps, and $c_{1},\ldots,c_{k}>0$ real numbers that satisfy $\sum_{i=1}^{k}c_{i}n_{i}=n$. We say that the set $\{(B_{i},c_{i})\}_{i=1,\ldots,k}$ forms a Brascamp–Lieb datum.
Theorem 14 (Brascamp–Lieb inequality).
Assume that $\{(B_{i},c_{i})\}_{i=1,\ldots k}$ forms a Brascamp–Lieb datum. For non-negative $f_{i}\in L_{1}(H_{i})$, we have
$\int_{\mathbb{R}^{n}}\prod_{i=1}^{k}f_{i}(B_{i}x)^{c_{i}}\,dx\leq C\prod_{i=1}^{k}\left(\int_{H_{i}}f_{i}\right)^{c_{i}},$ (16)
where $C$ is determined by choosing centered Gaussians $f_{i}(x)=e^{-\langle A_{i}x,x\rangle}$ with $A_{i}$ positive definite.
Remark 15.
The Geometric Brascamp–Lieb inequality is a special case of (16). We note that (16) reduces to Hölder’s inequality if $H_{1}=\ldots=H_{k}=\mathbb{R}^{n}$ and each $B_{i}=I_{n}$, and hence $C=1$ and $\sum_{i=1}^{k}c_{i}=1$ in that case.
We say that two sets of Brascamp–Lieb data $\{(B_{i},c_{i})\}_{i=1,\ldots,k}$ and $\{(B^{\prime}_{i},c^{\prime}_{i})\}_{i=1,\ldots,k^{\prime}}$ are equivalent if $k^{\prime}=k$, $c^{\prime}_{i}=c_{i}$, $i=1,\ldots,k$, and there exists a linear isomorphism $\Phi_{i}:H_{i}\to H^{\prime}_{i}$ for $i=1,\ldots,k$ such that $B^{\prime}_{i}=\Phi_{i}\circ B_{i}$. It was proved by Carlen, Lieb, Loss [25] in the rank one case, and by Bennett, Carbery, Christ, Tao [17] in general that there exists a set of extremizers $f_{1},\ldots,f_{k}$ for (16) if and only if the Brascamp–Lieb datum $\{(B_{i},c_{i})\}_{i=1,\ldots,k}$ is equivalent to some Geometric Brascamp–Lieb datum. Therefore, Valdimarsson’s Theorem 9 provides a full characterization of the equality case in Theorem 14, as well.
The following reverse version of the Brascamp–Lieb inequality was proved by Barthe in [7] in the rank one case, and in [8] in general (see Barthe, Cordero-Erausquin [10] for a heat equation, and Lehec [42] for a probabilistic argument).
Theorem 16 (Barthe’s Reverse Brascamp–Lieb inequality).
Assume that $\{(B_{i},c_{i})\}_{i=1,\ldots,k}$ forms a Brascamp–Lieb datum. For non-negative $f_{i}\in L_{1}(H_{i})$, we have
$\int_{\mathbb{R}^{n}}^{*}\sup_{x=\sum_{i=1}^{k}c_{i}B_{i}^{*}x_{i},\,x_{i}\in H_{i}}\;\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,dx\geq D\prod_{i=1}^{k}\left(\int_{H_{i}}f_{i}\right)^{c_{i}},$ (17)
where $D$ is determined by choosing centered Gaussians $f_{i}(x)=e^{-\langle A_{i}x,x\rangle}$ with $A_{i}$ positive definite.
Remark 17.
The Geometric Barthe’s inequality is a special case of (17). We note that (17) is the Prékopa–Leindler inequality if $H_{1}=\ldots=H_{k}=\mathbb{R}^{n}$ and each $B_{i}=I_{n}$, and hence $D=1$ and $\sum_{i=1}^{k}c_{i}=1$ in that case.
Concerning extremals in Theorem 16, Lehec [42] proved that if there exist some Gaussian extremizers for Barthe’s Inequality (17), then the corresponding Brascamp–Lieb datum $\{(B_{i},c_{i})\}_{i=1,\ldots,k}$ is equivalent to some Geometric Brascamp–Lieb datum; therefore, the equality case of (17) can be understood via Theorem 12 in that case.
However, it is still not known whether having any extremizers in Barthe’s Inequality (17) yields the existence of Gaussian extremizers. One possible approach is to use iterated convolutions and renormalizations as in Bennett, Carbery, Christ, Tao [17] in the case of Brascamp–Lieb inequality.
There are three main methods of proof that work for proving both the Brascamp–Lieb inequality and its reverse form due to Barthe. In the paper [8] Barthe used optimal transportation to prove his Reverse Brascamp–Lieb inequality and reprove the Brascamp–Lieb inequality simultaneously. A heat equation argument was provided in the rank one case by Carlen, Lieb, Loss [25] for the Brascamp–Lieb Inequality and by Barthe, Cordero-Erausquin [10] for Barthe’s Reverse Brascamp–Lieb inequality. The general versions of both inequalities are proved via the heat equation approach by Barthe, Huet [12]. Finally, simultaneous probabilistic arguments for the two inequalities are due to Lehec [42].
We note that Chen, Dafnis, Paouris [27] and Courtade, Liu [28], as well, deal systematically with finiteness conditions in Brascamp–Lieb and Barthe’s inequalities. The importance of the Brascamp–Lieb inequality is shown by the fact that besides harmonic analysis, probability and convex geometry, it has been also even applied in number theory, see e.g. Guo, Zhang [40]. Various versions of the Brascamp–Lieb inequality and its reverse form have been obtained by Balogh, Kristály [6] Barthe [9], Barthe, Cordero-Erausquin [10], Barthe, Cordero-Erausquin, Ledoux, Maurey [11], Barthe, Wolff [13, 14], Bennett, Bez, Flock, Lee [15], Bennett, Bez, Buschenhenke, Cowling, Flock [16], Bobkov, Colesanti, Fragalà [19], Bueno, Pivarov [24], Chen, Dafnis, Paouris [27], Courtade, Liu [28], Duncan [31], Ghilli, Salani [39], Kolesnikov, Milman [41], Livshyts [44, 45], Lutwak, Yang, Zhang [49, 50], Maldague [51], Marsiglietti [52], Rossi, Salani [54, 55].
3. Proof of Theorem 6
In order to prove Theorem 6, we recall the proof of Theorem 5, see Liakopoulos [43]*Section 3. We will use the following three simple observations. For a convex body $M\subset\mathbb{R}^{n}$ with $o\in\operatorname{int}M$, the value of the gauge function $\|x\|_{M}$ on a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{M}=\inf\{\lambda\geq 0\colon x\in\lambda M\}$. Then, since $|M|=\frac{1}{n}\int_{S^{n-1}}\|u\|^{-n}_{M}\,{\rm d}u$, it holds that
$\int_{\mathbb{R}^{n}}e^{-\|x\|_{M}}\,{\rm d}x=\int_{0}^{\infty}e^{-r}nr^{n-1}|M|\,{\rm d}r=n!|M|,$ (18)
see Schneider [56]*(1.54) on p. 58.
We often use the fact that for a convex body $M\subset\mathbb{R}^{n}$ with $o\in\operatorname{int}M$ and any linear subspace $F\subset\mathbb{R}^{n}$, we have $\|x\|_{M}=\|x\|_{M\cap F}$ for $x\in F$. This yields that if $\oplus_{j=1}^{m}G_{j}=\mathbb{R}^{n}$ holds for the pairwise orthogonal linear subspaces $G_{1},\ldots,G_{m}\subset\mathbb{R}^{n}$, and $M_{j}\subset G_{j}$ is a $\dim G_{j}$-dimensional compact convex set with $o\in\operatorname{relint}M_{j}$ for $j=1,\ldots,m$, then $M=\operatorname{conv}\{M_{1},\ldots,M_{m}\}$ satisfies
$\|x\|_{M}=\sum_{i=1}^{m}\|P_{G_{j}}x\|_{M_{j}}$ (19)
for any $x\in\mathbb{R}^{n}$.
If there exists any independent subspace associated to $E_{1},\ldots,E_{k}$, then let these be $F_{1},\ldots,F_{\ell}\subset\mathbb{R}^{n}$. If there exists no independent subspace, then let $\ell=1$ and $F_{1}=\{o\}$. Next, let $\widetilde{F}=\oplus_{j=1}^{\ell}F_{j}$, and hence $F_{0}=\widetilde{F}^{\bot}$ is the dependent subspace associated with $E_{1},\ldots,E_{k}$.
We define
$f(x)=e^{-\|x\|_{K}},$ (20)
which is a log-concave function with $f(0)=1$, and it satisfies (cf. (18)),
$\int_{\mathbb{R}^{n}}f(y)^{n}\,{\rm d}y=\int_{\mathbb{R}^{n}}e^{-n\|y\|_{K}}\,{\rm d}y=\int_{\mathbb{R}^{n}}e^{-\|y\|_{\frac{1}{n}K}}\,{\rm d}y=n!\left|\frac{1}{n}K\right|=\frac{n!}{n^{n}}\cdot|K|.$ (21)
We claim that
$n^{n}\int_{\mathbb{R}^{n}}f(y)^{n}\,{\rm d}y\geq\prod_{i=1}^{k}\Big{(}\int_{E_{i}}f(x_{i})\,{\rm d}x_{i}\Big{)}^{c_{i}}.$ (22)
Equating the traces on the two sides of (9), we deduce that $d_{i}=\dim E_{i}$ satisfy
$\sum_{i=1}^{k}\frac{c_{i}d_{i}}{n}=1.$ (23)
For $z=\sum_{i=1}^{k}c_{i}x_{i}$ with $x_{i}\in E_{i}$, the log-concavity of $f$ and its definition (20) imply
$f\left(\frac{z}{n}\right)=f\left(\sum_{i=1}^{k}\frac{c_{i}d_{i}}{n}\cdot\frac{x_{i}}{d_{i}}\right)\geq\prod_{i=1}^{k}f\left(\frac{x_{i}}{d_{i}}\right)^{\frac{c_{i}d_{i}}{n}}=\prod_{i=1}^{k}f(x_{i})^{\frac{c_{i}}{n}}.$ (24)
Now, the monotonicity of the integral and Barthe’s inequality with $f_{i}(x)=f(x)$ for $x\in E_{i}$ and $i=1,\ldots,k$ yield that
[…]
Making the change of variable $y=\frac{z}{n}$, we conclude (22). Computing the right-hand side of (22), we have
$\int_{E_{i}}f_{i}(x_{i})\,{\rm d}x_{i}=\int_{E_{i}}e^{-\|x_{i}\|_{K}}\,{\rm d}x_{i}=\int_{E_{i}}e^{-\|x_{i}\|_{K\cap E_{i}}}\,{\rm d}x_{i}=d_{i}!|K\cap E_{i}|$ (26)
for $i=1,\ldots,k$. Therefore, (21), (22) and (26) yield (8).
Now, assume that equality holds in (8), and hence equality holds in the second inequality of (25), where we applied Barthe’s inequality. Since each $f_{i}$ is log-concave, $i=1,\ldots,k$, and $\operatorname{supp}f_{i}=E_{i}$, it follows from Theorem 12 and Lemma 13 that
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle}\prod_{F_{j}\subset E_{i}}h_{j}\left(P_{F_{j}}(x)\right)$ (27)
for $x\in E_{i}$, where
  • (i)
    $\theta_{i}>0$,
  • (ii)
    $h_{j}\in L_{1}(F_{j})$ is log-concave for $j=1,\ldots,\ell$, and hence bounded,
  • (iii)
    $A\colon F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ is a positive definite matrix such that the eigenspaces of $A$ are critical subspaces.
Since each $f_{i}$ has its unique maximum at $o$ which equals $1$, we may assume that each $h_{j}$ has its unique maximum at $o$ which is also equal to $1$. It follows that $h_{j}=e^{-\varphi_{j}}$ for an even convex function $\varphi_{j}:F_{j}\to[0,\infty)$ with $\varphi_{j}(o)=0$, $j=1,\ldots,\ell$, and hence for any $i=1,\ldots,k$, (27) reads as
$f_{i}(x)=e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle}\prod_{F_{j}\subset E_{i}}\exp\left(-\varphi_{j}\left(P_{F_{j}}(x)\right)\right),$
and any $x\in E_{i}$ satisfies that
$\|x\|_{K}=\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle+\sum_{F_{j}\subset E_{i}}\varphi_{j}\left(P_{F_{j}}(x)\right).$ (28)
Next, we claim that
$F_{\mathrm{dep}}=\{o\}.$ (29)
We suppose that there exists $\dim F_{\mathrm{dep}}\geq 1$, and seek a contradiction. As any $E_{i}$ can be written as
$E_{i}=\left(\oplus_{F_{j}\subset E_{i}}F_{j}\right)\oplus(F_{\mathrm{dep}}\cap E_{i}),$
it follows from (7) that
$\sum_{i=1}^{k}c_{i}P_{E_{i}\cap F_{\mathrm{dep}}}=\left.I_{n}\right|_{F_{\mathrm{dep}}}.$ (30)
Therefore, there exists some $E_{i}$ such that $E_{i}\cap F_{\mathrm{dep}}\neq\{o\}$, and let $x_{0}\in E_{i}\cap F_{\mathrm{dep}}\neq\{o\}$. We deduce from (28) and $P_{F_{j}}(x_{0})=o$ if $F_{j}\subset E_{i}$ that
$0=\lim_{t\to\infty}\frac{\|tx_{0}\|_{K}}{t^{2}}=\lim_{t\to\infty}\frac{\langle A(tx_{0}),(tx_{0})\rangle}{t^{2}}=\langle Ax_{0},x_{0}\rangle>0,$
which contradiction proves (29).
In turn, we deduce from (29), the definition of $F_{\rm dep}$ and (28) that
$\oplus_{j=1}^{\ell}F_{j}=$ $\mathbb{R}^{n},$ (31)
$\|x\|_{K}=$ $\sum_{F_{j}\subset E_{i}}\varphi_{j}\left(P_{F_{j}}(x)\right)\mbox{ for any $x\in E_{i}$.}$ (32)
Now, (31) proves part (i) of Theorem 6.
In order to verify (ii), notice that if $x\in F_{m}$ for an $m\in\{1,\ldots,\ell\}$, then $\varphi_{j}\left(P_{F_{j}}(x)\right)=\varphi_{j}(o)=0$ for $j\neq m$, and hence (32) yields that $\varphi_{m}(x)=\|x\|_{K}$. We deduce again from (32) that if $x\in E_{i}$, then
$\|x\|_{K}=\sum_{F_{j}\subset E_{i}}\left\|P_{F_{j}}(x)\right\|_{K}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(x)\right\|_{K},$ (33)
where we used that either $F_{j}\subset E_{i}$ or $E_{i}\subset F_{j}^{\bot}$ for any $j\in\{1,\ldots,\ell\}$.
Let
$M=\operatorname{conv}\{K\cap F_{j}\}_{j=1,\ldots\ell}.$
It follows from (19) that if $z\in\mathbb{R}^{n}$, then
$\|z\|_{M}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(z)\right\|_{K}.$ (34)
Since $M\subset K$, we deduce that
$\|z\|_{K}\leq\|z\|_{M}$ (35)
for any $z\in\mathbb{R}^{n}$. In order to prove $\|z\|_{K}\geq\|z\|_{M}$, we note that equality in (8) yields equality in the first inequality in (25) by the argument above. It follows that if $z\in\mathbb{R}^{n}$, then
$\left(e^{-\|\frac{z}{n}\|_{K}}\right)^{n}=\sup_{z=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\prod_{i=1}^{k}e^{-c_{i}\|x_{i}\|_{K}};$
or, equivalently,
$\|z\|_{K}=\inf_{z=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\sum_{i=1}^{k}c_{i}\|x_{i}\|_{K}=\inf_{z=\sum_{i=1}^{k}y_{i},\,y_{i}\in E_{i}}\sum_{i=1}^{k}\|y_{i}\|_{K}.$ (36)
Since we may assume that $\sum_{i=1}^{k}\|y_{i}\|_{K}<\|z\|_{K}+1$ in (36), and hence $\|y_{i}\|_{K}<\|z\|_{K}+1$ for $i=1,\ldots,k$, we deduce, by compactness, the existence of $y_{i}\in E_{i}$ for $i=1,\ldots,k$ such that
$z=\sum_{i=1}^{k}y_{i}\mbox{ \ and \ }\|z\|_{K}=\sum_{i=1}^{k}\|y_{i}\|_{K}.$ (37)
Now, apply (37), (33), the triangle inequality, then (37) again and, finally, (34) to obtain that
$\|z\|_{K}=$ $\sum_{i=1}^{k}\|y_{i}\|_{K}=\sum_{i=1}^{k}\sum_{j=1}^{\ell}\left\|P_{F_{j}}(y_{i})\right\|_{K}=\sum_{j=1}^{\ell}\sum_{i=1}^{k}\left\|P_{F_{j}}(y_{i})\right\|_{K}$
$\geq$ $\sum_{j=1}^{\ell}\left\|\sum_{i=1}^{k}P_{F_{j}}(y_{i})\right\|_{K}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(z)\right\|_{K}=\left\|z\right\|_{M}.$
Therefore, $\|z\|_{K}=\|z\|_{M}$ for any $z\in\mathbb{R}^{n}$ by (35), which proves part (ii), thus completing the proof of Theorem 6.
Acknowledgements Károly J. Böröczky was partially supported by Hungarian NKFIH grant no. 150613.
The research of Ferenc Fodor was supported by project no. 150151, which has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.