### v2

# Equality in Liakopoulos’s generalized dual Loomis–Whitney inequality via Barthe’s Reverse Brascamp–Lieb inequality

Abstract.
We use the characterization of the case of equality in Barthe’s Geometric Reverse Brascamp–Lieb inequality to characterize equality in Liakopoulos’s volume estimate in terms of sections by certain lower-dimensional linear subspaces.
1. Introduction and results
Geometric inequalities, in which the volume of a shape is given and we search for the extremal value of a quantity, have been considered since the time of the ancient Greeks and have been at the center of attention in Geometry and Analysis for the last two centuries. The most classical one is probably the Isoperimetric inequality. Let $B^{n}\subset\mathbb{R}^{n}$ be the origin-centered Euclidean unit ball, and let $X\subset\mathbb{R}^{n}$ be a compact set such that ${\rm int}\,X\neq\emptyset$ and its boundary $\partial X$ is an (n1)(n-1)-dimensional Lipschitz manifold. Let $|X|$ denote the Lebesgue measure of $X$. We define the surface area $S(X)$ of $X$ as $S(X)=\lim_{r\to 0^{+}}\frac{|X+rB^{n}|-|X|}{r}$. The Isoperimetric inequality states that
$S(X)\geq n|B^{n}|^{\frac{1}{n}}\cdot|X|^{\frac{n-1}{n}},$
with equality if and only if $X$ is a Euclidean ball. One of the major results of Bent Fuglede was that he proved the optimal stability version of the Isoperimetric inequality for star-shaped $X\subset\mathbb{R}^{n}$ with Lipschitz boundary in terms of the Hausdorff distance in a series of papers [33, 34, 35] in 1986–1991 (see Figalli, Maggi, Pratelli [32] and Fusco, Maggi, Pratelli [36] for optimal stability results in terms of volume difference). Here, the Isoperimetric inequality is a consequence of the Brunn–Minkowski inequality
$|\alpha X+\beta Y|^{\frac{1}{n}}\geq\alpha|X|^{\frac{1}{n}}+\beta|Y|^{\frac{1}{n}}$
for compact sets $X,Y\subset\mathbb{R}^{n}$ with $|X|,|Y|>0$ and $\alpha,\beta>0$, where equality holds if and only if $X$ and $Y$ are convex and homothetic; namely, $Y=\lambda X+z$ for some $\lambda>0$ and $z\in\mathbb{R}^{n}$. See Figalli, Maggi, Pratelli [32] for an elegant proof of the Brunn–Minkowski inequality including equality conditions and even a stability version based on optimal transport. In turn, the Brunn–Minkowski inequality has a functional form that is called the Prékopa–Leindler inequality, which we recall in a form most suitable for use in this paper: If $\sum_{i=1}^{k}c_{i}=1$ holds for $c_{1},\ldots,c_{k}>0$ with $k\geq 2$, then for any non-negative $f_{i}\in L_{1}(\mathbb{R}^{n})$, $i=1,\ldots,k$, we have
$\int_{\mathbb{R}^{n}}^{*}\sup_{x=\sum_{i=1}^{k}c_{i}x_{i}}\;\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,dx\geq\prod_{i=1}^{k}\left(\int_{\mathbb{R}^{n}}f_{i}\right)^{c_{i}},$ (1)
where $\int_{\mathbb{R}^{n}}^{*}$ stands for the outer integral, as the integrand on the left hand side may not be measurable (see Gardner [37] for a survey on various aspects of the Brunn–Minkowski inequality and the Prékopa–Leindler inequality). If $\int_{\mathbb{R}^{n}}f_{i}>0$ for all $i=1,,\ldots,k$, then equality in the Prékopa–Leindler inequality (1) implies that there exist $a_{1},\ldots,a_{k}>0$, $w_{1},\ldots,w_{k}\in\mathbb{R}^{n}$ and a log-concave function $\varphi$ such that $\sum_{i=1}^{k}c_{i}w_{i}=o$ and $f_{i}(x)=a_{i}\varphi(x+w_{i})$ for a.e. $x\in\mathbb{R}^{n}$ (see Dubuc [29] and Balogh, Kristály [6] for the equality case, and Figalli, van Hintum, and Tiba [31] for a stability version). Here, a function $f:\,\mathbb{R}^{n}\to[0,\infty)$ is log-concave if $f((1-\lambda)x+\lambda y)\geq f(x)^{1-\lambda}f(y)^{\lambda}$ holds for $x,y\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$; or equivalently, $f=e^{-\varphi}$ for a convex function $\varphi:\,\mathbb{R}^{n}\to(-\infty,\infty]$. If, in addition, $0<\int_{\mathbb{R}^{n}}f<\infty$, then $\operatorname{supp}f=\operatorname{cl}\{f>0\}$ is convex, $f$ is continuous on $\operatorname{int}\operatorname{supp}f$, $f$ is bounded, and for any $t$ with $0<t<\sup f$, the open set $\{f>t\}$ is convex and bounded.
In this note, we characterize equality in Liakopoulos’s inequality (Theorem 5) as a first step towards a possible stability version. However, the story of this paper originates from the classical Loomis–Whitney inequality (2) proved in [46].
Let $e_{1},\ldots,e_{n}$ denote an orthonormal basis of $\mathbb{R}^{n}$. For a compact set $K\subset\mathbb{R}^{n}$ with ${\rm dim}\,{\rm aff}\,K=m$, we write $|K|$ for the $m$-dimensional Lebesgue measure of $K$. For a proper linear subspace $E$ of $\mathbb{R}^{n}$ ($E\neq\mathbb{R}^{n}$ and $E\neq\{0\}$), let $P_{E}$ denote the orthogonal projection onto $E$.
Theorem 1 (Loomis, Whitney).
If $K\subset\mathbb{R}^{n}$ is compact, ${\rm aff}\,K=\mathbb{R}^{n}$ and $|K|>0$, then
$|K|^{n-1}\leq\prod_{i=1}^{n}|P_{e_{i}^{\bot}}K|,$ (2)
with equality if and only if $K=\oplus_{i=1}^{n}K_{i}$, where $\mathrm{aff}\,K_{i}$ is a line parallel to $e_{i}$.
In order to consider a generalization of the Loomis–Whitney inequality, we set $n]:=\{1,\ldots,n\}$, and for a non-empty proper subset $\sigma\subset[n]$, we define $E_{\sigma}=\operatorname{lin}\{e_{i}\}_{i\in\sigma}$. For $s\geq 1$, we say that the not necessarily distinct proper non-empty subsets $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ if each $j\in[n]$ is contained in exactly $s$ of $\sigma_{1},\ldots,\sigma_{k}$. The Bollobás–Thomason inequality [16] is as follows.
Theorem 2 (Bollobás, Thomason).
If $K\subset\mathbb{R}^{n}$ is compact, ${\rm aff}\,K=\mathbb{R}^{n}$ and $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ for an integer $s\geq 1$, then
$|K|^{s}\leq\prod_{i=1}^{k}|P_{E_{\sigma_{i}}}K|.$ (3)
Remarks.
  • (i)
[…]
  • (ii)
    Bennett, Carbery, Christ and Tao introduced the notion of a ”Geometric Brascamp–Lieb datum” in [14]. A collection of linear subspaces $E_{1},\ldots,E_{k}\subset\mathbb{R}^{n}$, each of dimension at least one, together with positive constants $c_{1},\ldots,c_{k}>0$ forms a Geometric Brascamp–Lieb datum if
    $\sum_{i=1}^{k}c_{i}P_{E_{i}}=I_{n},$ (4)
    where $I_{n}$ is the identity map on $\mathbb{R}^{n}$. The term geometric refers to the fact that the surjective linear maps $P_{E_{i}}\colon\mathbb{R}^{n}\to E_{i}$, $i=1,\ldots,k$ are orthogonal projections.
    Now, the subspaces $E_{\sigma_{i}}:={\rm lin}\{e_{j}:j\in\sigma_{i}\}$ in Theorem 2 satisfy
    $\sum_{i=1}^{k}\frac{1}{s}\,P_{E_{\sigma_{i}}}=I_{n};$ (5)
    that is, they form a Geometric Brascamp–Lieb datum, where each coefficient is $\frac{1}{s}$.
The equality case of the Bollobás–Thomason inequality (Theorem 2), based on Valdimarsson’s characterization theorem [55], has been known by experts for some time without a published proof. Let $s\geq 1$, and let $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ be an $s$-uniform cover of $[n]$. We say that the disjoint union $n]=\tilde{\sigma}_{1}\sqcup\ldots\sqcup\tilde{\sigma}_{\ell}$ forms the partition of $[n]$ induced by the $s$-uniform cover $\sigma_{1},\ldots,\sigma_{k}$ if $\{\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{\ell}\}$ consists of all non-empty distinct subsets of $[n]$ of the form $\cap_{i=1}^{k}\sigma_{i,0}$, where for each $i$, either $\sigma_{i,0}=\sigma_{i}$ or $\sigma_{i,0}=[n]\setminus\sigma_{i}$.
Theorem 3 (Folklore).
Let $K\subset\mathbb{R}^{n}$ be compact with ${\rm aff}\,K=\mathbb{R}^{n}$ and $|K|>0$, and let $\sigma_{1},\ldots,\sigma_{k}\subset[n]$ form an $s$-uniform cover of $[n]$ for an integer $s\in\{1,\ldots,n-1\}$. Then equality holds in (3) if and only if $K=\oplus_{i=1}^{\ell}P_{E_{\tilde{\sigma}_{i}}}K$, where $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{\ell}$ form the partition of $[n]$ induced by $\sigma_{1},\ldots,\sigma_{k}$.
We note that Ellis, Friedgut, Kindler, and Yehudayoff [28] even proved a stability version of the Bollobás–Thomason inequality Theorem 2 in the case relevant to information theory, and hence that of the Loomis–Whitney inequality as well.
The real starting point of our note is Meyer’s dual Loomis–Whitney inequality in [51], where equality holds for affine cross-polytopes.
Theorem 4 (Meyer).
If $K\subset\mathbb{R}^{n}$ is a compact convex set with $o\in\mathrm{int}\,K$, then
$|K|^{n-1}\geq\frac{n!}{n^{n}}\prod_{i=1}^{n}\left|K\cap e_{i}^{\bot}\right|,$ (6)
with equality if and only if $K=\operatorname{conv}\{\pm\lambda_{i}e_{i}\}_{i=1}^{n}$ for $\lambda_{i}>0$, $i=1,\ldots,n$.
Let us observe that, unlike in the case of the Loomis-Whitney inequality, the assumption that $K$ is convex is necessary in Meyer’s inequality (6).
We note that various reverse and dual Loomis–Whitney type inequalities were proved by Campi, Gardner, Gronchi [23], Brazitikos et al. [20, 21], Alonso-Gutiérrez et al. [1, 2]. In particular, according to (5), the following theorem, due to Liakopoulos [42], is a generalization of Meyer’s inequality on one hand, and is the generalized dual form of the Bollobás–Thomason inequality (Theorem 2) on the other hand.
Theorem 5 (Liakopoulos).
If $K\subset\mathbb{R}^{n}$ is a compact convex set with $o\in\operatorname{int}K$, and the linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$, with ${\rm dim}\,E_{i}=n_{i}\in\{1,\ldots,n-1\}$, form a Geometric Brascamp–Lieb datum; that is,
$\sum_{i=1}^{k}c_{i}P_{E_{i}}=I_{n},$ (7)
then
$|K|\geq\frac{\prod_{i=1}^{k}(n_{i}!)^{c_{i}}}{n!}\cdot\prod_{i=1}^{k}\left|K\cap E_{i}\right|^{c_{i}}.$ (8)
Like in Meyer’s inequality, the convexity of $K$ is a necessary condition in Liakopoulos’ inequality.
Given the linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ as part of a Geometric Brascamp–Lieb datum in Theorem 5, we say that $F=\cap_{i=1}^{k}E_{i,0}$ is an independent subspace if either $E_{i,0}=E_{i}$ or $E_{i,0}=E_{i}^{\bot}$ holds for any $i=1,\ldots,k$, and ${\rm dim}\,F\geq 1$ (see also the description above Theorem 9). In general, there may not exist any independent subspace associated with $E_{1},\ldots,E_{k}$; however, if $F\neq F^{\prime}$ are independent subspaces, then $F^{\prime}\subset F^{\bot}$.
Our main result characterizes equality in Liakopoulos’s Theorem 5.
Theorem 6.
We have equality in (8) in Theorem 5 if and only if
  • (i)
    $\oplus_{j=1}^{\ell}F_{j}=\mathbb{R}^{n}$ holds for the independent subspaces $F_{1},\ldots,F_{\ell}\subset\mathbb{R}^{n}$ associated with $E_{1},\ldots,E_{k}$, and
  • (ii)
    $K=\operatorname{conv}\{K\cap F_{j}\}_{j=1}^{\ell}$.
The structure of the paper is as follows. We introduce the Brascamp–Lieb inequality and Barthe’s reverse form, which play a crucial role in the proof of Theorem 6 in Section 2. We prove Theorem 6 in Section 3.
2. Equality in Barthe’s Geometric Reverse Brascamp–Lieb inequality and some related results
Let the linear subspaces $E_{1},\ldots,E_{k}\subset\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$ form a Geometric Brascamp–Lieb datum. If $\dim E_{i}=n_{i}$ for $i=1,\ldots,n$, then
$\sum_{i=1}^{k}c_{i}n_{i}=n$ (9)
by comparing traces in (4).
The Geometric Brascamp–Lieb inequality (Theorem 7) traces back to the works of Brascamp and Lieb [19] and Ball [3, 4] in the rank-one case, where $\dim E_{i}=1$ for every $i=1,\ldots,k$, and to Lieb [43], Barthe [7] and Carbery, Christ, and Tao [14] in the general setting. We note that a rank one Geometric Brascamp–Lieb datum is called a Parseval frame in computer science (cf. Casazza, Tran, and Tremain [25]).
Theorem 7 (Brascamp, Lieb, Ball, Barthe).
For any non-trivial linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and constants $c_{1},\ldots,c_{k}>0$ satisfying (4), and for non-negative $f_{i}\in L_{1}(E_{i})$, we have
$\int_{\mathbb{R}^{n}}\prod_{i=1}^{k}f_{i}(P_{E_{i}}x)^{c_{i}}\,dx\leq\prod_{i=1}^{k}\left(\int_{E_{i}}f_{i}\right)^{c_{i}}$ (10)
Remark 8.
When $E_{1}=\cdots=E_{k}=\mathbb{R}^{n}$, then the statement (10) becomes Hölder’s inequality, and $\sum_{i=1}^{k}c_{i}=1$ holds in that case (cf. (9)).
According to (5), the Geometric Brascamp–Lieb inequality (Theorem 7) directly yields the Bollobás–Thomason inequality (Theorem 5) by choosing $E_{i}=E_{\sigma_{i}}$, $c_{i}=\frac{1}{s}$ and $f_{i}=\mathbf{1}_{P_{E_{\sigma_{i}}}}$ for $i=1,\ldots,k$. Equality in Theorem 7 occurs, for example, when each $f_{i}$ is a centered Gaussian density like $f_{i}(x)=e^{-\pi|x|^{2}}$ for all $i=1,\ldots,k$, where $|\cdot|$ denotes the Euclidean norm on $\mathbb{R}^{n}$.
Partial progress on the equality case of the Geometric Brascamp–Lieb inequality was made by Barthe [7], Carlen, Lieb, and Loss [24], and Bennett, Carbery, Christ, and Tao [14]. The full characterization was later achieved by Valdimarsson [55]. Before stating Valdimarsson’s result, we establish the necessary notation. Assume that the subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and constants $c_{1},\ldots,c_{k}>0$ form a Geometric Brascamp–Lieb datum. We call a nontrivial linear subspace $V$ critical if
$\sum_{i=1}^{k}c_{i}\dim(E_{i}\cap V)=\dim V.$
According to [14], this criterion is equivalent to the decomposition
$E_{i}=(E_{i}\cap V)+(E_{i}\cap V^{\perp})\quad\text{for each }i=1,\ldots,k.$
In order to characterize equality in Theorem 5, we also recall the notions of independent and dependent subspaces associated with a Geometric Brascamp–Lieb datum. These notions were originally defined by Valdimarsson [55]. Let $J$ denote the set of all functions from $\{1,\ldots,k\}$ to $\{0,1\}$. For $\varepsilon\in J$, let $F_{(\varepsilon)}=\cap_{i=1}^{k}E_{i}^{(\varepsilon(i))}$, where $E_{i}^{(0)}=E_{i}$ and $E_{i}^{(1)}=E_{i}^{\bot}$ for $i=1,\ldots,k$. If $\dim F_{(\varepsilon)}\geq 1$, then we call $F_{(\varepsilon)}$ independent. Note that if $F$ and $F^{\prime}$ are two different independent subspaces, then $F^{\prime}\subset F^{\perp}$. Let $J_{0}=\{j\in J\colon\dim F_{(\varepsilon)}\geq 1\}$. We call the orthogonal complement $F_{\mathrm{dep}}$ of $\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}$ in $\mathbb{R}^{n}$ the dependent subspace. In this way, $\mathbb{R}^{n}$ decomposes into a direct sum of pairwise orthogonal linear subspaces
$\mathbb{R}^{n}=\left(\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}\right)\oplus F_{\mathrm{dep}}.$ (11)
Observe that there may not exist any independent subspaces, in which case $\mathbb{R}^{n}=F_{\mathrm{dep}}$. It may also happen that the dependent subspace is trivial; then $\mathbb{R}^{n}=\oplus_{\varepsilon\in J_{0}}F_{(\varepsilon)}$. In general, all independent subspaces and the dependent subspace (if $F_{\mathrm{dep}}$ is non-trivial) are critical subspaces.
For a non-zero linear subspace $L\subset\mathbb{R}^{n}$, we say that the matrix representing a linear transformation $A:L\to L$ is positive definite if $A$ is symmetric (i.e., $A^{T}=A$) and $x^{T}Ax>0$ for all $x\in L\setminus\{0\}$.
Theorem 9 (Valdimarsson).
For any proper linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and real numbers $c_{1},\ldots,c_{k}>0$ satisfying (4), let us assume that equality holds in the Brascamp–Lieb inequality (10) for non-negative $f_{i}\in L_{1}(E_{i})$, $i=1,\ldots,k$. If $F_{\mathrm{dep}}\neq\mathbb{R}^{n}$, then let $F_{1},\ldots,F_{\ell}$ be the independent subspaces, and if $F_{\mathrm{dep}}=\mathbb{R}^{n}$, then let $\ell=1$ and $F_{1}=\{0\}$. Then there exist $b\in F_{\mathrm{dep}}$ and $\theta_{i}>0$ for $i=1,\ldots,k$, integrable non-negative $h_{j}:\,F_{j}\to[0,\infty)$ for $j=1,\ldots,\ell$, and a positive definite matrix $A:F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ such that the eigenspaces of $A$ are critical subspaces and
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x-b\rangle}\prod_{F_{j}\subset E_{i}}h_{j}(P_{F_{j}}(x))\mbox{ \ \ \ for Lebesgue a.e. $x\in E_{i}$}.$ (12)
On the other hand, if for any $i=1,\ldots,k$, $f_{i}$ is of the form as in (12), then equality holds in (10) for $f_{1},\ldots,f_{k}$.
The subspaces $F_{i}$, $i=1,\ldots,\ell$ are called independent to express the fact that the functions $h_{j}$ defined on $F_{j}$ may be selected arbitrarily and independently of one another in Theorem 9.
Barthe [7] established a reverse version of the Geometric Brascamp–Lieb inequality.
Theorem 10 (Barthe).
For any non-trivial linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and $c_{1},\ldots,c_{k}>0$ satisfying (4), and for non-negative $f_{i}\in L_{1}(E_{i})$, we have
$\int_{\mathbb{R}^{n}}^{*}\sup_{x=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\;\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,dx\geq\prod_{i=1}^{k}\left(\int_{E_{i}}f_{i}\right)^{c_{i}}.$ (13)
Remark 11.
If $E_{1}=\cdots=E_{k}=\mathbb{R}^{n}$, then Theorem 10 becomes the Prékopa–Leindler inequality (1), because $P_{E_{i}}=I_{n}$ and thus $\sum_{i=1}^{k}c_{i}=1$ (cf. (9)).
We note that Barthe’s Geometric Reverse Brascamp–Lieb inequality (Theorem 10) was used by Liakopoulos to prove the dual form of Bollobás–Thomason inequality (Theorem 5) by choosing $f_{i}=1_{K\cap E_{i}}$ for $i=1,\dots,k$. In Section 3, we follow this argument to exploit the equality case. Böröczky, Kalantzopoulos, Xi [18] proved the following characterization of equality in Barthe’s inequality (13).
Theorem 12 ([18]).
For linear subspaces $E_{1},\ldots,E_{k}$ of $\mathbb{R}^{n}$ and $c_{1},\ldots,c_{k}>0$ satisfying (4), if $F_{\mathrm{dep}}\neq\mathbb{R}^{n}$, then let $F_{1},\ldots,F_{\ell}$ be the independent subspaces, and if $F_{\mathrm{dep}}=\mathbb{R}^{n}$, then let $\ell=1$ and $F_{1}=\{0\}$.
If equality holds in Barthe’s Geometric Reverse Brascamp–Lieb inequality (13) for non-negative $f_{i}\in L_{1}(E_{i})$ with $\int_{E_{i}}f_{i}>0$, $i=1,\ldots,k$, then
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x-b_{i}\rangle}\prod_{F_{j}\subset E_{i}}h_{j}\left(P_{F_{j}}(x-w_{i})\right)$ (14)
for Lebesgue a.e. $x\in E_{i}$, where
  • $\theta_{i}>0$, $b_{i}\in E_{i}\cap F_{\rm dep}$ and $w_{i}\in E_{i}$ for $i=1,\ldots,k$,
  • $h_{j}\in L_{1}(F_{j})$ is non-negative for $j=1,\ldots,\ell$, and, in addition, $h_{j}$ is log-concave if there exist $\alpha\neq\beta$ with $F_{j}\subset E_{\alpha}\cap E_{\beta}$,
  • $A:F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ is a positive definite matrix such that the eigenspaces of $A$ are critical subspaces.
On the other hand, if for any $i=1,\ldots,k$, $f_{i}$ is of the form as in (14) and equality holds for all $x\in E_{i}$ in (14), then equality holds in (13) for $f_{1},\ldots,f_{k}$.
We discuss some special cases of Theorem 12 that will be useful in the proof of Theorem 6. In Lemma 13 we use the notation of Theorem 12.
Lemma 13.
Let $f_{1},\ldots,f_{k}$ form a set of extremizers of Barthe’s Geometric Reverse Brascamp–Lieb inequality.
  • (i)
    If for each $\alpha=1,\ldots,k$, the remaining subspaces $\{E_{i}\}_{i\neq\alpha}$ span $\mathbb{R}^{n}$, then each $f_{i}$ is $\mathcal{H}^{n_{i}}$ a.e. equal to some log-concave function on $E_{i}$.
  • (ii)
    If each $f_{i}$ is log-concave, then one may assume that each $h_{j}$, $j=1,\ldots,\ell$, is log-concave, and hence (14) holds for all $x\in\mathrm{relint}\,\mathrm{supp}\,f_{i}$.
  • (iii)
    If each $f_{i}$ is even, then one may assume that each $b_{i}=w_{i}=o$ for $i=1,\ldots,k$, and each $h_{j}$ is even for $j=1,\ldots,\ell$ (and, in fact, each $h_{j}$ being even and log-concave can be assumed if each $f_{i}$ is even and log-concave).
Proof.
For (i), the condition that for any $\alpha=1,\ldots,k$, $\{E_{i}\}_{i\neq\alpha}$ spans $\mathbb{R}^{n}$ in Theorem 12 is equivalent to saying that for any $F_{j}$, $j=1,\ldots,\ell$, there exist $\alpha\neq\beta$ with $F_{j}\subset E_{\alpha}\cap E_{\beta}$.
In the argument for (ii) and (iii), if $F_{\mathrm{dep}}\neq\{o\}$, then we write $h_{0,i}(y)=e^{-\langle Ay,y-b_{i}\rangle}$ for $i=1,\ldots,k$ and $y\in E_{i}\cap F_{\mathrm{dep}}$, and if $F_{j}\subset E_{i}$ is an independent subspace, then we write $h_{j,i}\left(y\right)=h_{j}\left(y-P_{F_{j}}w_{i}\right)$ for $y\in F_{j}$.
For (ii), first we show that even if some independent subspace $F_{m}$ is contained in a single $E_{\alpha}$, then $h_{m}$ is Lebesgue a.e. equal to some log-concave function on $F_{m}$. In this case, let $z\in E_{\alpha}\cap F_{m}^{\bot}$ such that $h_{j,\alpha}(P_{F_{j}}z)>0$ whenever $F_{j}\subset E_{\alpha}$ and $j\neq m$, and for Lebesgue a.e. $x\in z+F_{m}$, we have
$f_{\alpha}(x)=\theta_{\alpha}h_{0,\alpha}\left(P_{F_{\rm dep}}z\right)h_{m,\alpha}\left(P_{F_{m}}x\right)\prod_{F_{j}\subset E_{\alpha},\&\,j\neq m}h_{j,\alpha}\left(P_{F_{j}}z\right).$
As $f_{\alpha}$ is log-concave, $h_{m,\alpha}$ agrees with a log-concave function for Lebesgue a.e. $y\in F_{m}$, and hence the same holds for $h_{\alpha}$.
Now, as log-concave functions are continuous on their support, we may assume that each $h_{j}$ is actually log-concave, and hence (14) holds for all $x\in\operatorname{relint}(\operatorname{supp}f_{i})$ for all $i=1,\ldots,k$.
To show (iii), one replaces each $h_{j,i}$ with the function $y\mapsto\sqrt{h_{j,i}(y)\cdot h_{j,i}(-y)}$ that is log-concave if $h_{j,i}$ is log-concave. ∎
Part (i) of Lemma 13 states that if for every $E_{\alpha}$, the remaining subspaces $E_{i}$, $i\neq\alpha$ of the Brascamp–Lieb datum span $\mathbb{R}^{n}$, then the functions $h_{j}$ involved in the extremizers in Theorem 12 are necessarily log-concave. The reason behind this, as already explained in [18], is as follows. If $F_{j}$ is one of the independent subspaces, then the coefficients satisfy $\sum_{E_{i}\supset F_{j}}c_{i}=1$. When the functions $f_{1},\ldots,f_{k}$ have the form given in (14), then the equality case of Barthe’s inequality (13) takes the form
$\int^{*}_{F_{j}}\sup_{x=\sum_{E_{i}\supset F_{j}}c_{i}x_{i}\atop x_{i}\in F_{j}}h_{j}\Big(x_{i}-P_{F_{j}}w_{i}\Big)^{c_{i}}\,dx=\prod_{E_{i}\supset F_{j}}\left(\int_{F_{j}}h_{j}\Big(x-P_{F_{j}}w_{i}\Big)\,dx\right)^{c_{i}}\left(=\int_{F_{j}}h_{j}(x)\,dx\right).$
Consequently, if there exist two distinct indices $\alpha\neq\beta$ such that $F_{j}\subset E_{\alpha}\cap E_{\beta}$, then the equality conditions in the Prékopa–Leindler inequality (1) force each $h_{j}$ to be log-concave. In contrast, if there is some $\alpha\in{1,\ldots,k}$ for which $F_{j}\subset E_{\beta}^{\perp}$ for all $\beta\neq\alpha$, then no restriction is imposed on $h_{j}$, and in this case $c_{\alpha}=1$.
Finally, we note that the Brascamp–Lieb inequality and Barthe’s reverse form have more general versions proved, for example, via optimal transport by Barthe [7] or using the heat equation by Barthe, Cordero-Erausquin [9] (see, for example, [17] or [18] for a discussion what is known about the equality case), but our paper is only concerned with the Geometric versions. The Brascamp–Lieb inequality has been applied in various fields in mathematics, like probability, harmonic analysis and convex geometry, and even in number theory, for example, by Guo, Zhang [39]. In addition, several versions of the Brascamp–Lieb inequality and Barthe’s reverse form have been proved by Balogh, Kristály [6] Barthe [8], Barthe, Cordero-Erausquin [9], Barthe, Cordero-Erausquin, Ledoux, Maurey [10], Barthe, Wolff [11, 12], Bennett, Bez, Buschenhenke, Cowling, Flock [13], Bobkov, Colesanti, Fragalà [15], Bueno, Pivovarov [22], Chen, Dafnis, Paouris [26], Courtade, Liu [27], Duncan [30], Ghilli, Salani [38], Kolesnikov, Milman [40], Lehec [41], Livshyts [44, 45], Lutwak, Yang, Zhang [47, 48], Maldague [49], Marsiglietti [50], Rossi, Salani [52, 53].
3. Proof of Theorem 6
In order to prove Theorem 6, we recall the proof of Theorem 5; see Liakopoulos [42]*Section 3. We will use the following three simple observations. For a convex body $M\subset\mathbb{R}^{n}$ with $o\in\operatorname{int}M$, the value of the gauge function $\|x\|_{M}$ on a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{M}=\inf\{\lambda\geq 0\colon x\in\lambda M\}$. Then, since $|M|=\frac{1}{n}\int_{S^{n-1}}\|u\|^{-n}_{M}\,{\rm d}u$, it holds that
$\int_{\mathbb{R}^{n}}e^{-\|x\|_{M}}\,{\rm d}x=\int_{0}^{\infty}e^{-r}nr^{n-1}|M|\,{\rm d}r=n!|M|,$ (15)
see Schneider [54]*(1.54) on p. 58.
We often use the fact that for a convex body $M\subset\mathbb{R}^{n}$ with $o\in\operatorname{int}M$ and any linear subspace $F\subset\mathbb{R}^{n}$, we have $\|x\|_{M}=\|x\|_{M\cap F}$ for $x\in F$. This yields that if $\oplus_{j=1}^{m}G_{j}=\mathbb{R}^{n}$ holds for the pairwise orthogonal linear subspaces $G_{1},\ldots,G_{m}\subset\mathbb{R}^{n}$, and $M_{j}\subset G_{j}$ is a $\dim G_{j}$-dimensional compact convex set with $o\in\operatorname{relint}M_{j}$ for $j=1,\ldots,m$, then $M=\operatorname{conv}\{M_{1},\ldots,M_{m}\}$ satisfies
$\|x\|_{M}=\sum_{i=1}^{m}\|P_{G_{j}}x\|_{M_{j}}$ (16)
for any $x\in\mathbb{R}^{n}$.
If there exist any independent subspaces associated with $E_{1},\ldots,E_{k}$, then let these be $F_{1},\ldots,F_{\ell}\subset\mathbb{R}^{n}$. If there exists no independent subspace, then let $\ell=1$ and $F_{1}=\{o\}$. Next, let $\widetilde{F}=\oplus_{j=1}^{\ell}F_{j}$, and hence $F_{0}=\widetilde{F}^{\bot}$ is the dependent subspace associated with $E_{1},\ldots,E_{k}$.
We define
$f(x)=e^{-\|x\|_{K}},$ (17)
which is a log-concave function with $f(0)=1$, and it satisfies (cf. (15)),
$\int_{\mathbb{R}^{n}}f(y)^{n}\,{\rm d}y=\int_{\mathbb{R}^{n}}e^{-n\|y\|_{K}}\,{\rm d}y=\int_{\mathbb{R}^{n}}e^{-\|y\|_{\frac{1}{n}K}}\,{\rm d}y=n!\left|\frac{1}{n}K\right|=\frac{n!}{n^{n}}\cdot|K|.$ (18)
We claim that
$n^{n}\int_{\mathbb{R}^{n}}f(y)^{n}\,{\rm d}y\geq\prod_{i=1}^{k}\Big(\int_{E_{i}}f(x_{i})\,{\rm d}x_{i}\Big)^{c_{i}}.$ (19)
Equating the traces on the two sides of (4), we deduce that $n_{i}=\dim E_{i}$ satisfy
$\sum_{i=1}^{k}\frac{c_{i}n_{i}}{n}=1.$ (20)
For $z=\sum_{i=1}^{k}c_{i}x_{i}$ with $x_{i}\in E_{i}$, the log-concavity of $f$ and its definition (17) imply
$f\left(\frac{z}{n}\right)=f\left(\sum_{i=1}^{k}\frac{c_{i}n_{i}}{n}\cdot\frac{x_{i}}{n_{i}}\right)\geq\prod_{i=1}^{k}f\left(\frac{x_{i}}{n_{i}}\right)^{\frac{c_{i}n_{i}}{n}}=\prod_{i=1}^{k}f(x_{i})^{\frac{c_{i}}{n}}.$ (21)
Now, the monotonicity of the integral and Barthe’s inequality (13) with $f_{i}(x)=f(x)$ for $x\in E_{i}$ and $i=1,\ldots,k$ yield that
$\int_{\mathbb{R}^{n}}f\left(\frac{z}{n}\right)^{n}\,{\rm d}z\geq\int_{\mathbb{R}^{n}}^{*}\sup_{z=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\prod_{i=1}^{k}f_{i}(x_{i})^{c_{i}}\,{\rm d}z\geq\prod_{i=1}^{k}\Big(\int_{E_{i}}f_{i}(x_{i})\,dx_{i}\Big)^{c_{i}}.$ (22)
Making the change of variable $y=\frac{z}{n}$, we conclude (19). Computing the right-hand side of (19), we have
$\int_{E_{i}}f_{i}(x_{i})\,{\rm d}x_{i}=\int_{E_{i}}e^{-\|x_{i}\|_{K}}\,{\rm d}x_{i}=\int_{E_{i}}e^{-\|x_{i}\|_{K\cap E_{i}}}\,{\rm d}x_{i}=n_{i}!|K\cap E_{i}|$ (23)
by (15) for $i=1,\ldots,k$. Therefore, (18), (19) and (23) yield (8).
Now, assume that equality holds in (8), and hence equality holds in the second inequality of (22), where we applied Barthe’s inequality. Since each $f_{i}$ is log-concave, $i=1,\ldots,k$, and $\operatorname{supp}f_{i}=E_{i}$, it follows from Theorem 12 and Lemma 13 that
$f_{i}(x)=\theta_{i}e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle}\prod_{F_{j}\subset E_{i}}h_{j}\left(P_{F_{j}}(x)\right)$ (24)
for $x\in E_{i}$, where
  • (i)
    $\theta_{i}>0$,
  • (ii)
    $h_{j}\in L_{1}(F_{j})$ is log-concave for $j=1,\ldots,\ell$, and hence bounded,
  • (iii)
    $A\colon F_{\mathrm{dep}}\to F_{\mathrm{dep}}$ is a positive definite matrix such that the eigenspaces of $A$ are critical subspaces.
Since each $f_{i}$ has its unique maximum at $o$ which equals $1$, we may assume that each $h_{j}$ has its unique maximum at $o$ which is also equal to $1$. It follows that $h_{j}=e^{-\varphi_{j}}$ for an even convex function $\varphi_{j}:F_{j}\to[0,\infty)$ with $\varphi_{j}(o)=0$, $j=1,\ldots,\ell$, and hence for any $i=1,\ldots,k$, (24) reads as
$f_{i}(x)=e^{-\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle}\prod_{F_{j}\subset E_{i}}\exp\left(-\varphi_{j}\left(P_{F_{j}}(x)\right)\right),$
and any $x\in E_{i}$ satisfies that
$\|x\|_{K}=\langle AP_{F_{\mathrm{dep}}}x,P_{F_{\mathrm{dep}}}x\rangle+\sum_{F_{j}\subset E_{i}}\varphi_{j}\left(P_{F_{j}}(x)\right).$ (25)
Next, we claim that
$F_{\mathrm{dep}}=\{o\}.$ (26)
We suppose that there exists $\dim F_{\mathrm{dep}}\geq 1$, and seek a contradiction. As any $E_{i}$ can be written as
$E_{i}=\left(\oplus_{F_{j}\subset E_{i}}F_{j}\right)\oplus(F_{\mathrm{dep}}\cap E_{i}),$
it follows from (7) that
$\sum_{i=1}^{k}c_{i}P_{E_{i}\cap F_{\mathrm{dep}}}=\left.I_{n}\right|_{F_{\mathrm{dep}}}.$ (27)
Therefore, there exists some $E_{i}$ such that $E_{i}\cap F_{\mathrm{dep}}\neq\{o\}$, and let $x_{0}\in E_{i}\cap F_{\mathrm{dep}}\neq\{o\}$. We deduce from (25) and $P_{F_{j}}(x_{0})=o$ if $F_{j}\subset E_{i}$ that
$0=\lim_{t\to\infty}\frac{\|tx_{0}\|_{K}}{t^{2}}=\lim_{t\to\infty}\frac{\langle A(tx_{0}),(tx_{0})\rangle}{t^{2}}=\langle Ax_{0},x_{0}\rangle>0,$
and this contradiction proves (26).
In turn, we deduce from (26), the definition of $F_{\rm dep}$ and (25) that
$\oplus_{j=1}^{\ell}F_{j}=$ $\mathbb{R}^{n},$ (28)
$\|x\|_{K}=$ $\sum_{F_{j}\subset E_{i}}\varphi_{j}\left(P_{F_{j}}(x)\right)\mbox{ for any $x\in E_{i}$.}$ (29)
Now, (28) proves part (i) of Theorem 6.
In order to verify (ii), notice that if $x\in F_{m}$ for an $m\in\{1,\ldots,\ell\}$, then $\varphi_{j}\left(P_{F_{j}}(x)\right)=\varphi_{j}(o)=0$ for $j\neq m$, and hence (29) yields that $\varphi_{m}(x)=\|x\|_{K}$. We deduce again from (29) that if $x\in E_{i}$, then
$\|x\|_{K}=\sum_{F_{j}\subset E_{i}}\left\|P_{F_{j}}(x)\right\|_{K}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(x)\right\|_{K},$ (30)
where we used that either $F_{j}\subset E_{i}$ or $E_{i}\subset F_{j}^{\bot}$ for any $j\in\{1,\ldots,\ell\}$.
Let
$M=\operatorname{conv}\{K\cap F_{j}\}_{j=1,\ldots\ell}.$
It follows from (16) that if $z\in\mathbb{R}^{n}$, then
$\|z\|_{M}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(z)\right\|_{K}.$ (31)
Since $M\subset K$, we deduce that
$\|z\|_{K}\leq\|z\|_{M}$ (32)
for any $z\in\mathbb{R}^{n}$. In order to prove $\|z\|_{K}\geq\|z\|_{M}$, we note that equality in (8) yields equality in the first inequality in (22) by the argument above. It follows that if $z\in\mathbb{R}^{n}$, then
$\left(e^{-\|\frac{z}{n}\|_{K}}\right)^{n}=\sup_{z=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\prod_{i=1}^{k}e^{-c_{i}\|x_{i}\|_{K}};$
or, equivalently,
$\|z\|_{K}=\inf_{z=\sum_{i=1}^{k}c_{i}x_{i},\,x_{i}\in E_{i}}\sum_{i=1}^{k}c_{i}\|x_{i}\|_{K}=\inf_{z=\sum_{i=1}^{k}y_{i},\,y_{i}\in E_{i}}\sum_{i=1}^{k}\|y_{i}\|_{K}.$ (33)
Since we may assume that $\sum_{i=1}^{k}\|y_{i}\|_{K}<\|z\|_{K}+1$ in (33), and hence $\|y_{i}\|_{K}<\|z\|_{K}+1$ for $i=1,\ldots,k$, we deduce, by compactness, the existence of $y_{i}\in E_{i}$ for $i=1,\ldots,k$ such that
[…]
Now, apply (34), (30), the triangle inequality, then (34) again, and finally (31) to obtain that
$\|z\|_{K}=$ $\sum_{i=1}^{k}\|y_{i}\|_{K}=\sum_{i=1}^{k}\sum_{j=1}^{\ell}\left\|P_{F_{j}}(y_{i})\right\|_{K}=\sum_{j=1}^{\ell}\sum_{i=1}^{k}\left\|P_{F_{j}}(y_{i})\right\|_{K}$
$\geq$ $\sum_{j=1}^{\ell}\left\|\sum_{i=1}^{k}P_{F_{j}}(y_{i})\right\|_{K}=\sum_{j=1}^{\ell}\left\|P_{F_{j}}(z)\right\|_{K}=\left\|z\right\|_{M}.$
Therefore, $\|z\|_{K}=\|z\|_{M}$ for any $z\in\mathbb{R}^{n}$ by (32), which proves part (ii), thus completing the proof of Theorem 6.
Acknowledgments
We are grateful to the two reviewers whose remarks substantially improved the presentation.
Károly J. Böröczky was partially supported by Hungarian NKFIH grant no. 150613. The research of Ferenc Fodor was supported by NKFIH project no. 150151, which has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.